# Vivid-VR: Distilling Concepts from Text-to-Video Diffusion Transformer for Photorealistic Video Restoration

### v2

@@ 5.3 Effect of the Concept Distillation Strategy
To mitigate distribution drift caused by imperfect multimodal alignment in training data, we introduce a concept distillation training strategy, that leverages the T2V base model to generate training data. To verify the effectiveness of this strategy, we disable the generated training data and train this baseline method only on the collected videos. The results in Table 2(e) show that the baseline method without the concept distillation strategy fails to achieve high-quality results and the performance degrades significantly ($1.47dB$ lower in DOVER). The visual results in Figure 7(f) confirms this, showing overly sharp textures due to distribution drift. For the similar reason, textures generated by STAR [37], DOVE [8] and SeedVR [31] are also less realistic. In contrast, the proposed method adopting concept distillation strategy performs significantly better in terms of texture vividness and realism. In addition, as shown in Figure 5(f), the baseline method without the concept distillation strategy also suffers from a decline in temporal coherence. This verifies that our concept distillation method can facilitate video restoration in terms of both perceptual quality and temporal consistency.
@@ 6 Conclusions and Limitations
We have proposed Vivid-VR, a DiT-based generative video restoration method built upon an advanced T2V foundation model. To mitigate distribution drift during fine-tuning, we have introduced a concept distillation training strategy that leverages the pre-trained T2V model to synthesize training samples with embedded textual concepts. Regarding the model architecture for controllable generation, we have proposed two key components: 1) a control feature projector that removes degradation artifacts from latent video features, and 2) a dual-branch ControlNet connector combining an MLP and cross-attention mechanism for control feature mapping and dynamic retrieval. Both quantitative and qualitative experimental results demonstrate the effectiveness of the proposed method.

### v3

@@ 5.3 Effect of the Concept Distillation Strategy
To mitigate distribution drift caused by imperfect multimodal alignment in training data, we introduce a concept distillation strategy, that leverages the T2V base model to generate training data. To verify its effectiveness, we disable the generated training data and train this baseline method only on the collected videos. Table 2(f) and Figure 7(e) show that the baseline method without the concept distillation strategy fails to achieve high-quality results, showing overly sharp textures due to distribution drift. For the similar reason, textures generated by STAR and SeedVR are also less realistic. In addition, the baseline method without the concept distillation also suffers from a decline in temporal coherence (see Figure 5(f)). This verifies that our concept distillation can facilitate video restoration in terms of both perceptual quality and temporal consistency.
Furthermore, to verify whether using a more accurate VLM captioner could resolve the distribution drift problem, we conduct an ablation study: using the more advanced Qwen2.5-VL as VLM captioner for training data annotation. As evidenced by Table 2(g) and Figure 7(f), the more accurate VLM captioner (i.e., Qwen2.5-VL) also fails to completely eliminate the modality gap in the T2V model’s latent space, demonstrating the persistence of distribution drift even with superior captioning models.
@@ 6 Conclusions and Limitations
We have proposed Vivid-VR, a DiT-based generative video restoration method built upon an advanced T2V foundation model. To mitigate distribution drift during fine-tuning, we introduced a concept distillation strategy that leverages the pre-trained T2V model to synthesize training samples with embedded textual concepts. Regarding the model architecture for controllable generation, we proposed two key components: 1) a control feature projector for degradation artifacts removal, and 2) a dual-branch connector combining an MLP and cross-attention mechanism for control feature mapping and dynamic retrieval. Both quantitative and qualitative experimental results demonstrate the effectiveness of the proposed method.
@@
Appendix A Appendix
We present additional analysis of the proposed method here. First, we discuss the trade-off between fidelity and quality in the proposed Vivid-VR. Second, we further analyze the effect of the number of training videos generated by the proposed concept distillation strategy. Furthermore, we discuss the effect of the number of ControlNet blocks. Finally, we present more visualization results.
Trade-off between Fidelity and Quality. As [41] points out, powerful generative prior is a double-edged sword, as excessive generative capacity may in turn affect the fidelity of the restored video. To address this, we introduce Restoration-Guided Sampling into Vivid-VR’s inference sampling process to balance the quality and fidelity:
$\hat{x}^{est}_{t}=x^{est}_{t}+(\frac{t}{T})^{\tau}(x^{lq}-x^{est}_{t}),$ (3)
where $x^{est}$ is the denoised latent at time step $t$, and $x^{lq}$ is the original input latent; $T$ denotes the total number of denoising steps; $\tau$ is the guidance coefficient; $\hat{x}^{est}$ is the output latent after the restoration-guided sampling. Figure 8 demonstrates this trade-off: higher guidance coefficient $\tau$ yield more realistic results, while lower $\tau$ preserve greater fidelity to the original input content.
Effect of the Number of Generated Training Videos. As mentioned in the main paper, the proposed method employs the concept distillation strategy to generate $100K$ videos for training. A natural question arises: does the number of generated training videos impact restoration performance? To investigate this, we conducted the ablation studies here. Table 3 demonstrates that increasing the number of generated training videos from 0 to $100K$ yields significant performance gains, while expanding from $100K$ to $150K$ shows diminishing returns. Considering the cost of generating training videos, we therefore adopt $100K$ generated videos as our standard configuration. Furthermore, we verified that relying solely on generated training data (without source videos) leads to suboptimal results. This occurs because the T2V base model’s outputs contain inherent imperfections. Training exclusively on such data ultimately compromises model performance.
Table 3: Effect of the number of training videos generated by the proposed concept distillation.
Training Videos NIQE MUSIQ CLIP-IQA DOVER
# Source # Generated
(a) 500K 0 5.364 57.36 0.363 12.99
(b) 500K 50K 4.562 63.00 0.408 13.46
(c) 500K 100K 4.361 67.61 0.450 14.46
(d) 500K 150K 4.292 67.19 0.443 14.51
(e) 0 150K 5.652 53.77 0.377 11.63
Effect of the Number of ControlNet Blocks. To reduce the parameter count, we employ $N/7$ DiT blocks in ControlNet. In [9], only one block is used, and all connectors share the same control feature. To further investigate whether $N/7$ DiT blocks are indeed necessary, we set the number of blocks to $1$ and retrain using the same settings. The results in Table 4 show that using only one block does not perform well.
Table 4: Effect of the number of ControlNet blocks.
NIQE MUSIQ CLIP-IQA DOVER
11 block 4.855 66.85 0.442 14.17
$N/7$ blocks (Vivid-VR) 4.361 67.61 0.450 14.46
More Visualization Results. In the main paper, we have demonstrated that the proposed Vivid-VR achieves state-of-the-art performance. Furthermore, we provide more visual comparisons on the project homepage (https://csbhr.github.io/projects/vivid-vr/), where Vivid-VR demonstrates superior structural clarity, texture richness, and visual vividness.
Figure 8: Trade-off between fidelity and quality. Higher guidance coefficient $\tau$ in the Restoration-Guided Sampling yield more realistic results, while lower $\tau$ preserve greater fidelity to the original input content. (Zoom-in for best view)